You purchased 100 shares of Chipotle Mexican Grill, Inc. (CMG) common stock on margin at

$431.79 per share.  Assume the initial margin is 50% and the maintenance margin is 30%. One

year later, the stock price closes at $837.11. If the broker’s call loan rate is 2.00%, what is your

return on equity?
Answer #1

Equity = Initial margin = 431.79 * 100 * 50% = 21,589.50

Broker's Loan = Remaining 50% = 21589.50

Interest on call loan = 21589.50 * 2% = 431.79

Net Gain = Number of shares * ( Closing stock price - Purchase price ) - Interest on call loan

= 100 * ( 837.11 - 431.79 ) - 431.79

= 100 * 405.32 - 431.79

= 40100.21

Return on Equity = Net Gain / Equity * 100

= 40100.21 / 21,589.50 * 100

= 185.74 %